OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? (x − 2)2 + (y + 3)2 = 5 (x + 2)2 + (y − 3)2 = 25 (x − 2)2 + (y + 3)2 = 25 (x − 2)2 − (y + 3)2 = 5

OpenStudy (alekos):

standard form is (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the centre and r is the radius

OpenStudy (anonymous):

i dont understand how to do it

OpenStudy (anonymous):

Is it the third one?

OpenStudy (alekos):

centre is (2,-3) so h=2 and k = -3

OpenStudy (anonymous):

So the second one

OpenStudy (alekos):

if you substitute the h and k values you get (x-2)^2 +(y+3)^2 = r^2 so now you have to work out the radius r

OpenStudy (alekos):

i'll let you do that

OpenStudy (anonymous):

25

OpenStudy (anonymous):

So the second one

OpenStudy (alekos):

the radius is the distance between the two points, which is?